OpenStudy (anonymous):

Laplace transforms problem sin3t+cos3t

OpenStudy (amistre64):

have you got a table setup?

OpenStudy (amistre64):

otherwise its just a matter if integrateing a by parts with e^(-st)

OpenStudy (amistre64):

\[\large\int sin(at)~e^{-st}~dt\] sin(at) -e^(-st)/s a cos(at) dt e^(-st) dt \[\large -\frac{1}{s}e^{-st}~sin(at)+\frac{a}{s}\int cos(at)~e^{-st}~dt\] cos(at) -e^(-st)/s -a sin(at) dt e^(-st) dt etc ...

OpenStudy (amistre64):

at 0 and infinity i believe that first term goes zero .....

OpenStudy (anonymous):

someone help me with this please

OpenStudy (unklerhaukus):

\[\mathcal L\Big\{\sin (nt)\Big\}(s)=\frac{n}{s^2+n^2}\] \[\mathcal L\Big\{\cos (nt)\Big\}(s)=\frac{s}{s^2+n^2}\]

OpenStudy (amistre64):

please do not post your own question into someone elses post .... butyes, someone can help with that